# The Inadequacy of Similarity-based Privacy Metrics:

### v4

@@ 8.2.4 Limiting Number of Calls
Alternatively, the provider could set global or per-user limits on the number of API calls. However, this would also undermine the promise of unlimited data. Moreover, detecting an unusually high volume of calls in real time could be challenging, as generating large amounts of data to simulate various edge cases is often part of normal testing activities. More importantly, detection may be unfeasible altogether, as synthetic data platforms are frequently deployed on clients’ premises or private clouds [49, 124, 83, 57], where providers lack direct access to monitor and mitigate such risks. We also demonstrated that, in some settings, even a handful of calls can enable very accurate attacks, leaving to future work to optimize the number of calls to obtain optimal trade-offs. Incidentally, as discussed in Sec. 5.1, providers allow users to generate large numbers of queries, which is also a standard assumption in previous related attacks [16, 84].
@@ Appendix A Reconstruction of All Train Data
To validate the hypothesis that reconstructing outliers is inherently more difficult, we conduct an experiment aimed at recovering any train records from Adult. We impose stricter constraints than those in Sec. 6, i.e., we limit SampleAttack to only 250 rounds, down from 1,000, and allow SearchAttack to trace just 3 distances instead of 4. The results are summarized in Table 4.
We observe that even under these limited settings, ReconSyn reconstructs a significant proportion of the train data. Specifically, using only a fraction of the computations, the average recovery rate of any train data is 0.89%, which exceeds the average recovery rate of outliers for Adult (average of 0.85%) as presented in Sec. 6. Overall, this should not be a surprise, given that outliers are less likely to be generated. The rightmost column in Table 4 further supports this, showing a substantial drop in recovering the outliers for all models, with PrivBayes being the only exception.
@@ Appendix B Datasets
2d Gauss. We sample 2,000 points from a standard bivariate normal distribution with zero correlation. We consider all train points beyond the blue circle displayed in Fig. 4 (centered at 0, radius 2.15) outliers, or about 10%.
@@
Adult. We use two versions of the Adult dataset [30]. We randomly sample 6,000 data points and refer to this dataset as Adult. Then, we further simplify this dataset by selecting six columns (‘age,’ ‘education,’ ‘marital status,’ ‘relationship,’ ‘sex,’ and ‘income’), and denote the resulting dataset as Adult Small. For both, we fit a Gaussian Mixture model with 10 clusters; for the former, we select the smallest cluster to be outliers, while for the latter, the smallest two (2d UMAP reduction shown in Fig. 10(a) and 10(b)).
Census. We randomly sample 10,000 data points from the Census dataset [30]. To determine the outliers, we fit a Gaussian Mixture model with 4 clusters and select the smallest one (2d UMAP reduction is plotted in Fig. 10(c)).
@@
MNIST. We sample 9,000 data points from the digits 3, 5, 8, and 9 from the MNIST [71] dataset as well as 1,000 from 0 and treat them as outliers (2d UMAP reduction displayed in Fig. 10(d)). To simplify the dataset, we downscale the images to 8x8 pixels and discretize all pixels to 16 bins.
Outliers Definition. While various definitions of outliers exist in literature [13, 77], we define underrepresented data regions and outliers to capture various scenarios, aiming to intuitively select roughly 10% of the train data (to serve as targets, as noted in Table 5). For 2d Gauss, we identify outliers as points lying beyond a certain distance from the center. In Adult Small, Adult, and Census, outliers are the smallest clusters determined by a Gaussian Mixture model. For MNIST, the digit 0 is deliberately underrepresented. We apply the same strategy/model to the test data. UMAP’s distance-preserving feature allows these outlier identification strategies to be visually verified in Fig. 10(a)10(d).
@@ Appendix D Additional SBPMs Counter-Examples
CE4. Metrics Inconsistency. Suppose we rely on the oracle to sample 1,000 new datasets A good privacy metric should reflect that the oracle perfectly preserves the train data privacy by reporting a high privacy score. Only on 274 occasions (out of 1,000), all the privacy tests pass. This demonstrates that metrics and empirical approaches measuring the privacy of a single synthetic dataset completely fail to capture the generating process.
@@
CE5. OF & Metrics Inconsistency. We also examine how applying the OF, which is supposed to improve privacy, affects it according to the metrics. Again, we rely on the oracle to draw synthetic data samples. On the left plot of Fig. 11(a), we see that the synthetic data passes all 3 tests, while on the right plot, when the outliers are filtered out, both DCR and NNDR fail. Even more surprisingly, Fig. 11(b) shows that removing the outliers can cause a previously failing test to pass (NNDR) and vice versa (DCR). These examples further prove the untrustworthiness and inconsistency of the privacy metrics and filters. One could observe all possible combinations of DCR and NNDR passing/failing.
@@
CE6. Discretization Inconsistency. Finally, in Fig. 12, we measure the effect of discretizing data. We test two discretization strategies, uniform and quantile, while varying the number of bins ($n_{bins}$) from 2 to 1,000. For discrete data, we use Hamming distance, while for continuous data, Euclidean. We test on 25d Gauss, use an oracle to sample 1,000 synthetic datasets, and report averages.

### v5

@@ 8.2.4 Limiting Number of Calls
Alternatively, the provider could set global or per-user limits on the number of API calls. However, this would also undermine the promise of unlimited data. Moreover, detecting an unusually high volume of calls in real time could be challenging, as generating large amounts of data to simulate various edge cases is often part of normal testing activities. More importantly, detection may be unfeasible altogether, as synthetic data platforms are frequently deployed on clients’ premises or private clouds [49, 124, 83, 57], where providers lack direct access to monitor and mitigate such risks. We also demonstrated that, in some settings, even a handful of calls can enable very accurate attacks, leaving to future work to optimize the number of calls to obtain optimal trade-offs. Incidentally, as discussed in Section 5.1, providers allow users to generate large numbers of queries, which is also a standard assumption in previous related attacks [16, 84].
@@ Appendix A Reconstruction of All Train Data
To validate the hypothesis that reconstructing outliers is inherently more difficult, we conduct an experiment aimed at recovering any train records from Adult. We impose stricter constraints than those in Section 6, i.e., we limit SampleAttack to only 250 rounds, down from 1,000, and allow SearchAttack to trace just 3 distances instead of 4. The results are summarized in Table 4.
We observe that even under these limited settings, ReconSyn reconstructs a significant proportion of the train data. Specifically, using only a fraction of the computations, the average recovery rate of any train data is 0.89%, which exceeds the average recovery rate of outliers for Adult (average of 0.85%) as presented in Section 6. Overall, this should not be a surprise, given that outliers are less likely to be generated. The rightmost column in Table 4 further supports this, showing a substantial drop in recovering the outliers for all models, with PrivBayes being the only exception.
@@ Appendix B Datasets
2d Gauss. We sample 2,000 points from a standard bivariate normal distribution with zero correlation. We consider all train points beyond the blue circle displayed in Figure 4 (centered at 0, radius 2.15) outliers, or about 10%.
@@
Adult. We use two versions of the Adult dataset [30]. We randomly sample 6,000 data points and refer to this dataset as Adult. Then, we further simplify this dataset by selecting six columns (‘age,’ ‘education,’ ‘marital status,’ ‘relationship,’ ‘sex,’ and ‘income’), and denote the resulting dataset as Adult Small. For both, we fit a Gaussian Mixture model with 10 clusters; for the former, we select the smallest cluster to be outliers, while for the latter, the smallest two (2d UMAP reduction shown in Figure 10(a) and 10(b)).
Census. We randomly sample 10,000 data points from the Census dataset [30]. To determine the outliers, we fit a Gaussian Mixture model with 4 clusters and select the smallest one (2d UMAP reduction is plotted in Figure 10(c)).
@@
MNIST. We sample 9,000 data points from the digits 3, 5, 8, and 9 from the MNIST [71] dataset as well as 1,000 from 0 and treat them as outliers (2d UMAP reduction displayed in Figure 10(d)). To simplify the dataset, we downscale the images to 8x8 pixels and discretize all pixels to 16 bins.
Outliers Definition. While various definitions of outliers exist in literature [13, 77], we define underrepresented data regions and outliers to capture various scenarios, aiming to intuitively select roughly 10% of the train data (to serve as targets, as noted in Table 5). For 2d Gauss, we identify outliers as points lying beyond a certain distance from the center. In Adult Small, Adult, and Census, outliers are the smallest clusters determined by a Gaussian Mixture model. For MNIST, the digit 0 is deliberately underrepresented. We apply the same strategy/model to the test data. UMAP’s distance-preserving feature allows these outlier identification strategies to be visually verified in Figure 10(a)10(d).
@@ Appendix D Additional SBPMs Counter-Examples
CE4. Metrics Inconsistency. Suppose we rely on the oracle to sample 1,000 new datasets. A good privacy metric should reflect that the oracle perfectly preserves the train data privacy by reporting a high privacy score. Only on 274 occasions (out of 1,000), all the privacy tests pass. This demonstrates that metrics and empirical approaches measuring the privacy of a single synthetic dataset completely fail to capture the generating process.
@@
CE5. OF & Metrics Inconsistency. We also examine how applying the OF, which is supposed to improve privacy, affects it according to the metrics. Again, we rely on the oracle to draw synthetic data samples. On the left plot of Figure 11(a), we see that the synthetic data passes all 3 tests, while on the right plot, when the outliers are filtered out, both DCR and NNDR fail. Even more surprisingly, Figure 11(b) shows that removing the outliers can cause a previously failing test to pass (NNDR) and vice versa (DCR). These examples further prove the untrustworthiness and inconsistency of the privacy metrics and filters. One could observe all possible combinations of DCR and NNDR passing/failing.
@@
CE6. Discretization Inconsistency. Finally, in Figure 12, we measure the effect of discretizing data. We test two discretization strategies, uniform and quantile, while varying the number of bins ($n_{bins}$) from 2 to 1,000. For discrete data, we use Hamming distance, while for continuous data, Euclidean. We test on 25d Gauss, use an oracle to sample 1,000 synthetic datasets, and report averages.
@@
Regarding the randomly looking score drops, we have two hypotheses. First, most drops (except 4) appear when the number of bins is below 100. This could be because with larger numbers of bins, we discretize the space into exponentially many hypercubes ($n_{bins}^{25}$ since there are 25 columns), and it becomes increasingly unlikely to get two numbers in the same hypercube even if their actual Euclidean distance is small. Second, the most drops occur when using the uniform strategy (at least for IMS and DCR). This could be because of the nature of the dataset (i.e., normal distribution). For uniform discretization, naturally, the bins in the middle will contain most of the data records (while there will be very few in the tails), increasing the chance of a match.